OpenStudy (anonymous):

w/24 >(and equal to) -1/6

OpenStudy (anonymous):

\[\frac{w}{24} \ge \frac{-1}{6}\]multiply through by 24\[w \ge \frac{-24}{6}=-4\]in interval notation\[[-4, \infty)\]graphically